A population of values has a normal distribution with u = 91.9 and o = 38.4. A random sample of size n = 232 is drawn.


- What is the mean of the distribution of sample means? Ux=


-What is the standard deviation of the distribution of sample means? Round your answer to two decimal places. Ox=





1
Expert's answer
2022-05-05T17:00:03-0400

1.


"\mu_{\bar{x}}=\mu=91.9"

2.

"\sigma_{\bar{x}}=\dfrac{\sigma}{\sqrt{n}}=\dfrac{38.4}{\sqrt{232}}=2.52"